# All loop soft photon theorems and higher spin currents on the celestial sphere

### v2

@@ IX Algebra of higher spin currents
The algebra (90) subject to the restriction (89) is well known to be the wedge subalgebra $\wedge w_{1+\infty}$ Bakas:1989xu ; Pope:1991zka ; Pope:1991ig . This is a classical algebra because we have neglected all the terms involving more than one contraction. If $k=0$ then of course this reduces to an infinite dimensional Abelian algebra. In the context of tree level soft graviton theorems $w_{1+\infty}$ has already appeared as a symmetry of the dual Celestial CFT2 Strominger:2021mtt ; Adamo:2021lrv .
@@ IX.1 Conformal transformation of higher spin currents
X Modifications of currents at finite $k$ and their impact on single soft theorems
As we pointed out earlier, the current algebra involving $\bar{d}^{j}$ for $j\geq\frac{3}{2}$ does not close among themselves at the quantum level. For example, this is evident from the presence of the blue terms in (83) for the OPE of two spin-$\frac{3}{2}$ currents. Thus, at the quantum level, in order to demand a closed algebra among currents of spin $j\geq\frac{3}{2}$, we need to modify the definitions of currents in (58) in the following way:
$\bar{D}^{\frac{3}{2}}{\alpha_{1}\alpha_{2}\alpha_{3}}=\bar{d}^{\frac{3}{2}}_{\alpha_{1}\alpha_{2}\alpha_{3}}+kA\left(\epsilon_{\alpha_{1}\alpha_{2}}\bar{\partial}\bar{d}^{\frac{1}{2}}_{\alpha_{3}}+\text{symmetrization}\right),$ (100)
$\begin{gathered}\bar{D}^{2}{\alpha_{1}\alpha_{2}\alpha_{3}\alpha_{4}}=\bar{d}^{2}_{\alpha_{1}\alpha_{2}\alpha_{3}\alpha_{4}}+kB\left(\epsilon_{\alpha_{1}\alpha_{2}}:\bar{d}^{\frac{1}{2}}_{\alpha_{3}}\bar{\partial}\bar{d}^{\frac{1}{2}}_{\alpha_{4}}:+\text{symmetrization}\right)\\ +kC\left(\epsilon_{\alpha_{1}\alpha_{2}}\bar{\partial}\bar{d}^{1}_{\alpha_{3}\alpha_{4}}+\text{symmetrization}\right)\end{gathered}$ (101)
where $A,B,C$ are dimensionless numbers independent of $k$ and are purely quantum corrections since $k\sim i\hbar$. Possibly they can be fixed by demanding that the algebra among $\bar{D}^{j}$ closes along the same line as in Bergshoeff:1991un . Now the question is whether, for non-vanishing $A,B,C$ the soft theorems receive modifications starting at the 3-loop order if we replace $\bar{d}^{j}$ by $\bar{D}^{j}$ in (71). The answer is no: the soft theorem Ward identities in (16) do not receive corrections as the $\bar{D}^{j}$s appear contracted with $2j$ numbers of $\bar{W}$ and $\bar{W}^{\alpha_{1}}\bar{W}^{\alpha_{2}}\cdots\bar{W}^{\alpha_{2j}}\bar{D}^{j}_{\alpha_{1}\alpha_{2}\cdots\alpha_{2j}}=\bar{W}^{\alpha_{1}}\bar{W}^{\alpha_{2}}\cdots\bar{W}^{\alpha_{2j}}\bar{d}^{j}_{\alpha_{1}\alpha_{2}\cdots\alpha_{2j}}$ even when $A,B,C\cdots$ above are non-vanishing. So there is no way to read off the values of $A,B,C\cdots$ from single soft photon theorem expressions.
XI What is the value of $k$?
@@ Acknowledgment
We would like to thank Alok Laddha, Ana-Maria Raclariu and Ashoke Sen for very helpful discussions. SB would also like to thank all the participants of the Pre-ISM Workshop in TIFR Mumbai, where this work was presented, for interesting questions and comments. We acknowledge the use of ChatGPT for providing general information and for assisting with exploratory checks to arrive at the projection operator (6). The work of SB is supported by the Swarnajayanti Fellowship (File No- SB/SJF/2021-22/14) of the Department of Science and Technology and ANRF, India. The work of RM is supported by ANRF-CRG project (CRG/2022/006165). The work of BS was supported by STFC under grant ST/X000753/1.
@@ Appendix A Support for the correctness of conjectured $n$-loop soft factor
Based on the soft photon theorem derivation at order $\omega(\ln\omega)^{2}$ in Sahoo:2020ryf from the analysis of two-loop Feynman diagrams in four dimensional scalar QED in presence of scalar contact interaction, we find that for scattering of massless charged particles the only diagrams giving non-vanishing contributions are those in which one end of both the virtual photon propagators are attached to the charged particle line from which the soft photon is emitted. Rest of the diagrams give vanishing contributions to $\mathcal{O}\left(\omega(\ln\omega)^{2}\right)$ soft factor, when charged particles are massless. [^11] Note that massless charged particles can themselves mediate long-range interactions and may generate additional contributions to logarithmic soft photon factors. This can be readily verified by replacing one or more virtual photon lines in Fig. 1 with massless charged scalar propagators, allowing for the corresponding three-point interaction vertices among them. However, if interaction vertices involving three massless charged particles are disallowed, one can argue using power counting that the coefficients $\mathbb{S}^{(n)}_{\text{em}}$ are not modified in the absence of any other massless particles besides the photon and charged particles. After examining a few higher-loop orders, this feature appears to be generic for massless charged particle scattering. Now assuming this observation to be valid for all-loop order S-matrix in massless scalar QED, we find that the only $n$-loop Feynman diagrams in Fig.1 provide non-vanishing contributions to $\mathbb{S}_{\text{em}}^{(n-1)}$, when evaluated with G-photon virtual propagators Sahoo:2018lxl ; Sahoo:2020ryf . The first diagram in Fig.1 represents the set of diagrams in which all the $n$ virtual photon (blue wavy lines) propagators are attached between particle-$a$ and $n$ different charged particles (black solid lines) through scalar–scalar–photon vertices, and the soft photon is attached to particle-$a$ at the far right through a scalar–scalar–photon vertex. The second diagram in Fig.1 represents the set of diagrams in which all but one of the $n$ virtual photon propagators are attached between particle-$a$ and $n$ different charged particles through scalar–scalar–photon vertices, while the soft photon is attached to particle-$a$ at the far right through a scalar–scalar–photon–photon vertex where one of the remaining virtual photon is also attached. The third diagram in Fig.1 represents the set of diagrams that differ from the first diagram only in that two or more virtual photons are now attached between particle-$a$ and particle-$b_{i}$ via ladder-type or cross-ladder-type configurations. Similarly, the fourth diagram in Fig.1 represents the set of diagrams that differ from the second diagram only in that two or more virtual photons are attached between particle-$a$ and particle-$b_{i}$ via ladder-type or cross-ladder-type configurations.
@@
From the ratio of two amplitudes as follows from (1), the leading-log $n$-loop soft factor from the set of the Feynman diagrams in Fig.1 turns out to be
@@
where $\Lambda$ is of order $\sqrt{|p_{b}.p_{c}|}$, and in the second line after using the sign $\simeq$ we only kept the $\mathcal{O}(\ln\omega)$ contribution. Note that, while evaluating the above integral following Sahoo:2018lxl , one encounters collinear divergences if the computation is performed directly in the massless case. However, all such collinearly divergent terms cancel once the sum over the particle indices $b,c$ is performed. Hence the massless limit of the integration result provided Sahoo:2018lxl for massive charged particles does not suffer from any collinear divergences. Following the terminology of Sahoo:2018lxl , we refer to the first term inside the parentheses, namely $\delta_{\eta_{b}\eta_{c},1}$, as the “classical” contribution, and the second term as the “quantum” contribution. For a careful analysis of the “classical” part of the above integral directly in the massless case, the reader is referred to the appendix of Choi:2025mzg . Now substituting the expression for $K_{\text{em}}$ in (104) we find the following soft factor coefficient at order $\omega^{n-1}(\ln\omega)^{n}$:

### v3

@@ IX Algebra of higher spin currents
The algebra (94) subject to the restriction (93) is well known to be the wedge subalgebra $\wedge w_{1+\infty}$ Bakas:1989xu ; Pope:1991zka ; Pope:1991ig . This is a classical algebra because we have neglected all the terms involving more than one contraction. If $k=0$ then of course this reduces to an infinite dimensional Abelian algebra. In the context of tree level soft graviton theorems $\wedge w_{1+\infty}$ has already appeared as a symmetry of the dual Celestial CFT2 Strominger:2021mtt ; Adamo:2021lrv .
@@ IX.1 Conformal transformation of higher spin currents
X What is the value of $k$?
@@ Acknowledgment
We are grateful to Andrew Strominger for pointing out a conceptual problem in the previous version of this paper. We are also grateful to Alok Laddha, Andrea Puhm, Ana-Maria Raclariu and Ashoke Sen for very helpful discussions. SB would also like to thank all the participants of the Pre-ISM Workshop in TIFR Mumbai, where this work was presented, for interesting questions and comments. We acknowledge the use of ChatGPT for providing general information and for assisting with exploratory checks to arrive at the projection operator (9). The work of SB is supported by the Swarnajayanti Fellowship (File No- SB/SJF/2021-22/14) of the Department of Science and Technology and ANRF, India. The work of RM is supported by ANRF-CRG project (CRG/2022/006165). The work of BS was supported by STFC under grant ST/X000753/1.
@@ Appendix A Support for the correctness of conjectured $n$-loop soft factor
Building on the derivation of the $\mathcal{O}(\omega(\ln\omega)^{2})$ soft photon theorem in four-dimensional massive scalar QED with scalar contact interactions, obtained from an explicit analysis of two-loop Feynman diagrams in Sahoo:2020ryf , we revisit the result at the leading order in the high-energy limit given in (3). We find that, after performing the loop integrals and subsequently taking the high-energy limit (3), the only non-vanishing contributions at the leading order arise from diagrams in which one end of each of the two virtual photon propagators is attached to the same charged particle line that emits the soft photon. All remaining two-loop diagrams, some of which do contribute to the $\mathcal{O}(\omega(\ln\omega)^{2})$ soft factor at general kinematics of charged particle scattering does not contribute to the leading order in the high-energy expansion (3). After examining a few higher-loop orders, this feature appears to be generic for leading order contribution to soft photon factor in the high-energy limit (3). Now assuming this observation to be valid for all-loop order S-matrix in scalar QED, we find that the only $n$-loop Feynman diagrams in Fig.1 provide non-vanishing contributions to $\mathbb{S}_{\text{em}}^{(n-1)}$ at leading order in the high-energy limit (3), when evaluated with G-photon virtual propagators Sahoo:2018lxl ; Sahoo:2020ryf . The first diagram in Fig.1 represents the set of diagrams in which all the $n$ virtual photon (blue wavy lines) propagators are attached between particle-$a$ and $n$ different charged particles (black solid lines) through scalar–scalar–photon vertices, and the soft photon is attached to particle-$a$ at the far right through a scalar–scalar–photon vertex. The second diagram in Fig.1 represents the set of diagrams in which all but one of the $n$ virtual photon propagators are attached between particle-$a$ and $n$ different charged particles through scalar–scalar–photon vertices, while the soft photon is attached to particle-$a$ at the far right through a scalar–scalar–photon–photon vertex where one of the remaining virtual photon is also attached. The third diagram in Fig.1 represents the set of diagrams that differ from the first diagram only in that two or more virtual photons are now attached between particle-$a$ and particle-$b_{i}$ via ladder-type or cross-ladder-type configurations. Similarly, the fourth diagram in Fig.1 represents the set of diagrams that differ from the second diagram only in that two or more virtual photons are attached between particle-$a$ and particle-$b_{i}$ via ladder-type or cross-ladder-type configurations.
@@
From the ratio of two amplitudes as follows from (1), the leading-log $n$-loop soft factor at leading order in the high-energy limit (3) from the set of the Feynman diagrams in Fig.1 turns out to be
@@
where $\Lambda$ is of order $\sqrt{|p_{b}.p_{c}|}$, and in the second line after using the sign $\simeq$ we only kept the $\mathcal{O}(\ln\omega)$ contribution. Following the terminology of Sahoo:2018lxl , we refer to the first term inside the parentheses, namely $\delta_{\eta_{b}\eta_{c},1}$, as the “classical” contribution, and the second term as the “quantum” contribution. [^12] For a careful analysis of the “classical” part of the above integral in the massless case, the reader is referred to the appendix of Choi:2025mzg . Now substituting the expression for $K_{\text{em}}$ in (106) we find the following soft factor coefficient at order $\omega^{n-1}(\ln\omega)^{n}$ and at the leading order in the high-energy limit (3):